OpenStudy (anonymous):

A two-digit number is 7 times the sum of its digits. The tens digit is 3 more than the units digit. What is the number?

OpenStudy (anonymous):

So I'm thinking our first equation will be: xy=7(x + y)

OpenStudy (anonymous):

10x+y=7(x+y) x=y+3.Now solving them

OpenStudy (diyadiya):

Oh wait

OpenStudy (anonymous):

The number is not xy @algebraprobs it is 10x+y

OpenStudy (anonymous):

Oh......oh yeah because it is in the tens place!

OpenStudy (diyadiya):

Yes @pavankumartgpk is absolutely Right Sorry about that!@algebraprobs

OpenStudy (anonymous):

Now you also have x=y+3 right.So substituting x=2y here gives 2y=y+3 so y=3 and x=6

OpenStudy (anonymous):

Number is 63

OpenStudy (anonymous):

Thanks! : )